# IR Bounds on Theories with Spontaneously-Broken Lorentz Symmetry

### v2

@@ 2.1 Scalar gapless mode
When solved perturbatively in $c_{2}$ and $c_{3}$, this gives,
@@
As expected, due to the conformal symmetry in $D=3$ and the breaking of Lorentz invariance below $\mu$, the speed of propagation at zero momentum is $c_{s}^{2}=1/2$. In general we find a momentum dependent phase velocity [^2] In this case the usual definition holds since the background is spatially isotropic and there is no gap. A more general definition is provided in Sec. 3. $\omega_{\pi}^{2}({\bf p})/{\bf p}^{2}$, which either increases or decreases with respect to $1/2$ depending on the sign of $c_{2}+c_{3}$. At this point, we only notice that the phase velocity will increase monotonically for increasing (small) momenta only if $c_{2}+c_{3}>0$. However, there is no further bound we can impose at this stage on the size of these coefficients, other than the requirement of validity of the perturbative treatment within the EFT, which is rather uninformative.
2.2 Dynamical gauge field
As discussed before, in general for finite $b$ and $d$ the quadratic Lagrangian (6) describes not only a propagating scalar, but also a propagating gauge field. In fact, these are kinetically mixed and therefore the eigenmodes of propagation are a combination of both. Interestingly, this endows the would-be scalar mode with a gap.
@@ 2.2 Dynamical gauge field
where $\Phi_{I}=\left(\bar{\pi},\bar{A}^{\nu}\right)$, and $\bar{\pi}=\sqrt{\mu}\,\pi$ and $\bar{A}^{\nu}=\frac{1}{\sqrt{\mu}}A^{\nu}$ are the dimension-$1/2$ fields. The kinetic $4\times 4$ matrix $\mathcal{K}_{IJ}(\omega,{\bf p})$ is given by
$\mathcal{K}_{IJ}(\omega,{\bf p})=\begin{pmatrix}c_{1}G^{-1}_{\pi\pi}&i\hat{J}_{0}&i\hat{J}_{j}\\ -i\hat{J}_{0}&u&v\,p_{j}\\ -i\hat{J}_{i}&v\,p_{i}&V\delta_{ij}+Wp_{i}p_{j}\\ \end{pmatrix}\,,$ (10)
with
@@
00 0 $=$ $V\,,$ (17)
00 0 $=$ $u\left(V+W{\bf p}^{2}\right)-v^{2}{\bf p}^{2}\,.$ (18)
The first equation describes the transverse mode, with dispersion relation
@@
Since there are no contributions from the higher-dimensional operators $c_{2}$ and $c_{3}$ in $\omega_{\perp}^{2}({\bf p})$, we will not examine this further. The second one is a mix of the scalar and $A^{0}$ modes (after removing the longitudinal mode by the gauge fixing condition). In this second dispersion relation the size of $c_{2}$ and $c_{3}$ controls contributions with higher powers in $\omega$ and/or ${\bf p}$. Due to this, Eq. (18) contains in fact both low-energy and high-energy modes. Focusing on the low-energy mode, we can solve perturbatively at linear order in $c_{2,3}$ within the regime of validity of the EFT. This leads to:
@@
Notice that when $c_{1}\ll(b+d)$, the vector effectively decouples from the scalar and also $m^{2}\to 0$, thus recovering the gapless scalar mode from Eq. (8). Keeping instead $b\,,\,d$ finite, the above equation has the form:
@@
One would be tempted to use the coefficient of the ${\bf p}^{2}$ term in Eq. (22), namely $\alpha^{2}({\bf p})$, as a notion of velocity. However, as we will discuss in the next section, this requires some care due to the presence of a gap. After clarifying this, we will use the dispersion relation (22) to derive bounds in Sec. 4.
3 Ray velocity from the acoustic metric
In order to study the propagation of the low-energy modes in this theory, it is useful to turn to the effective description given by the acoustic metric. The purpose of this Section is to provide a general definition of velocity that can be used for gapped modes and arbitrary backgrounds, and which reduces to the usual phase velocity for gapless modes. This is the so-called ray velocity Boillat:1969slt ; Boillat:1970gw ; Landau , which is given in Eq. (32), and which will prove important for the formulation of our bounds. The reader interested only on the bounds may skip this section and instead go directly to Sec. 4. For a more detailed account of this formalism see Ref. Sawicki:2024ryt .
We begin by considering a theory with gapless excitations described by a quadratic Lagrangian of the form
@@ 3 Ray velocity from the acoustic metric
where $Z^{\mu\nu}$ is the acoustic metric and $S_{\mu\nu}$ its inverse, and $S=\det(S_{\mu\nu})$. In this case the associated linearized equation of motion (EoM) can be written as
@@
with ¯$\bar{\square}=Z^{\mu\nu}\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}$ and ¯$\bar{\nabla}_{\mu}$ is the covariant derivative compatible with $Z$ and $S$. This is simply the wave-equation on the nontrivial background given by the acoustic metric.
It is possible to study this equation in the so called Eikonal expansion, with an ansatz $\delta\chi=A(x)\,e^{i\mathcal{S}(x)/\epsilon}$, and $\epsilon\to 0$. This amounts to consider a rapidly-varying phase $\mathcal{S}(x)$ and a slowly-varying amplitude $A(x)$. One can show that at leading order in $1/\epsilon$, that is the geometric optics approximation, Eq. (25) corresponds to
$Z^{\mu\nu}p_{\mu}p_{\nu}=0\,,$ (26)
with $p_{\mu}=\partial_{\mu}\mathcal{S}$, which effectively picks up the terms in the linearized EoM with the most derivatives acting on the fluctuations. To properly regard $Z^{\mu\nu}$ as an effective metric, we should be able from the above to derive a corresponding geodesic equation. Acting with the covariant derivative ¯$\bar{\nabla}_{\alpha}$ on the above we get
@@
where we used the fact that $p_{\mu}$ is the gradient of a scalar quantity, the phase $\mathcal{S}$, and therefore ¯$\bar{\nabla}_{\alpha}p_{\mu}=\bar{\nabla}_{\mu}p_{\alpha}$. Finally, contracting with $Z^{\nu\alpha}$ and defining the “ray-vector” as
$N^{\mu}\equiv Z^{\mu\nu}p_{\nu}\,,$ (28)
we obtain the geodesic equation
@@
In light of this, Eq. (27) can be rephrased as the parallel transport of energy-momentum along the $N^{\mu}$ geodesic, $N^{\mu}\bar{\nabla}_{\mu}p_{\nu}=0$. Physically, the ray vector $N^{\mu}$ tracks the flow of energy-momentum $p_{\mu}$. It is also useful to use the definition of $N^{\mu}$ to rewrite the EoM as
@@
In conclusion, the gapless EoM implies that the contravariant vector $N^{\mu}$ follows null geodesics of $S_{\mu\nu}$. This defines the trajectories, a.k.a “rays”, describing propagation of the wavefronts in the geometric optics approximation,
$N^{\mu}p_{\mu}=N^{\mu}\partial_{\mu}\mathcal{S}=0\quad,\quad\text{for gapless modes.}$. (31)
Since $N^{\mu}$ is the tangent vector of the rays, it is natural to define the speed of propagation along the geodesic – the ray velocity – in a given coordinate frame as
${\bf v}=\frac{{\bf N}}{N^{0}}\,.$ (32)
This is a general definition which is valid for arbitrary backgrounds, and can even accommodate the presence of a gap, as we will explain below in Sec. 3.2. In the case of a gapless mode, $N^{\mu}$ parametrizes the trajectories of the constant phase surfaces as expressed by Eq. (31), meaning that Eq. (32) coincides with the phase velocity of a gapless mode Landau ; Sawicki:2024ryt . Note that $N^{\mu}$ should be normalized to have dimensions of a length.
3.1 An example: P(X) theory
To see that this treatment reproduces the usual known results, let us turn to the simple example of a $P(X)$ theory with Lagrangian
The corresponding acoustic metric can be obtained by expanding at quadratic order in perturbations around a fixed background, $\phi=\bar{\phi}+\varphi$, and it reads
which is gapless. We proceed to compute the inverse of $Z^{\mu\nu}$, which is easily found to be
which allows to write the EoM in the geometric optics approximation in terms of $N^{\mu}$, obtaining
From here we can compute the ray velocity as defined in Eq. (32),
which coincides with the usual definition of phase velocity $\omega^{2}({\bf p})/{\bf p}^{2}$ for a background with spatially isotropic acoustic metric [^3] See the discussion in Sawicki:2024ryt . If we impose $N^{\mu}p_{\mu}=0$, then it follows that $\frac{{\bf N}\cdot{\bf p}}{N^{0}\omega}=1$, meaning that $\omega^{2}/{\bf p}^{2}={\bf N}^{2}/(N^{0})^{2}$, provided ${\bf p}$ is parallel to ${\bf N}$..
Suppose that in this theory we want to avoid superluminality. Requiring the ray velocity (38) to be less or equal to the speed of light $c=1$ immediately implies [^4] Here we are assuming that the denominator above is always positive, which is guaranteed by the absence-of-ghosts condition.
$c_{2}\geq 0\,.$ (39)
This is the well known condition to prevent superluminality in this model Adams:2006sv . Geometrically, this means the ray-vector must be timelike (or null) according to the standard spacetime metric $\eta_{\mu\nu}$, i.e. lie inside (or on) the lightcone. Using Eq. (37) above
which again implies (39).
As a closing remark, in this case due to the absence of a mass term, we could have also obtained this result by instead demanding that the momentum vector lies outside (or on) the lightcone, that is, it is $\eta$-spacelike (or null). From the EoM in terms of $p_{\mu}$ we get
again leading to (39). However, in general this does not imply that the ray velocity agrees with $\omega^{2}({\bf p})/{\bf p}^{2}$. This is especially important when comparing it against a $c_{s}^{2}\neq 1$ or when handling a theory with a mass gap.
3.2 Beyond the zero-gap approximation
For our particular purpose of studying the gapped mode (22) in the low-momentum regime, we must consider the opposite limit where momenta are not necessarily large enough to allow us to neglect the terms in the EoM with less derivatives. To keep the discussion simple we consider just a constant term [^5] This is justified in our case since we have a constant acoustic metric ($\partial_{\rho}Z^{\mu\nu}=0$, and therefore we can simply consider a plane-wave ansatz $\delta\chi=A\,e^{ip_{\mu}x^{\mu}}$, with $A$ and $p_{\mu}$ constants, instead of the more general Eikonal ansatz.,

### v3

@@ II.1 Scalar gapless mode
When solved perturbatively in $c_{2}$ and $c_{3}$ at linear order, this gives
@@
As expected, due to the conformal symmetry in $D=3$ and the breaking of Lorentz invariance below $\mu$, the speed of propagation at zero momentum is $c_{s}^{2}=1/2$. In general we find a momentum dependent phase velocity [^2] In this case the usual definition holds since the background is spatially isotropic. A more general definition is provided in Ref. [40]. $\omega_{\pi}^{2}({\bf p})/{\bf p}^{2}$, which either increases or decreases with respect to $1/2$ depending on the sign of $c_{2}+c_{3}$. At this point, we only notice that the phase velocity will increase monotonically for increasing (small) momenta only if $c_{2}+c_{3}>0$. However, there is no further bound we can impose at this stage on the size of these coefficients, other than the requirement of validity of the perturbative treatment within the EFT, which is rather uninformative. Importantly, the coefficients $c_{2}$ and $c_{3}$ only appear here in the specific combination $c_{2}+c_{3}$, an indication that the information contained in the bound (7) cannot be captured by the scalar mode alone. This prompts us to consider the dynamical nature of the gauge field in the search for an accurate representation of the analyticity bound.
II.2 Dynamical gauge field
As discussed before, in general, for finite $b$ and $d$, the quadratic Lagrangian (6) describes not only a propagating scalar, but also a propagating gauge field. In fact, these are kinetically mixed, and therefore, the eigenmodes of propagation are a combination of both. Interestingly, this endows the would-be scalar mode with a gap.
@@ II.2 Dynamical gauge field
where $\Phi_{I}=\left(\bar{\pi},\bar{A}_{\nu}\right)$, and $\bar{\pi}=\sqrt{\mu}\,\pi$ and $\bar{A}_{\nu}=\frac{1}{\sqrt{\mu}}A_{\nu}$ are the dimension-$1/2$ fields, with the indices $I,J=\{\chi,0,1,2\}$ running over the fields and their components. The kinetic $4\times 4$ matrix $\mathcal{K}_{IJ}(\omega,{\bf p})$ is given by
$\mathcal{K}_{IJ}(\omega,{\bf p})=\left(\begin{array}[]{c | c c }c_{1}G^{-1}_{\pi\pi}&i\hat{J}_{0}&i\hat{J}_{j}\\ \hline\cr-i\hat{J}_{0}&u&v\,p_{j}\\ -i\hat{J}_{i}&v\,p_{i}&V\delta_{ij}+b\,p_{i}p_{j}\\ \end{array}\right)\,$, (14)
which is organized into $1\times 1$, $1\times 3$, $3\times 1$ and $3\times 3$ blocks for clarity. The lower-case indices $i,j=\{1,2\}$ run instead over the spatial coordinates. The various quantities introduced here read as
@@
$0$ $=$ $V\,$, (20)
$0$ $=$ $u\left(V+b{\bf p}^{2}\right)-v^{2}{\bf p}^{2}\,.$ (21)
The first equation describes the transverse mode, with the dispersion relation
@@
Since there are no contributions from the HD operators $c_{2}$ and $c_{3}$ in $\omega_{\perp}^{2}({\bf p})$, we will not examine this further. The second one is a mix of the scalar and $A^{0}$ modes (after removing the longitudinal mode by the gauge fixing condition). In this second dispersion relation the size of $c_{2}$ and $c_{3}$ controls contributions with higher powers in $\omega$ and/or $\bf p$}. Due to this, Eq. (21) contains in fact both low-energy and high-energy modes. Focusing on the low-energy mode, we can solve perturbatively at linear order in $c_{2,3}$ within the regime of validity of the EFT. This leads to
@@
Notice that when $c_{1}\ll(b+d)$, the gauge effectively decouples from the scalar and also $m^{2}\ll\mu^{2}$, thus recovering the gapless scalar mode from Eq. (9). Keeping instead b,db\,,\,d finite, the above equation has the form:
@@
One would be tempted to use the coefficient of the ${\bf p}^{2}$ term in Eq. (25), namely $\alpha^{2}({\bf p})$, as a notion of velocity. However, as we will discuss in the next section, this requires some care due to the presence of a gap. After clarifying this, we will use the dispersion relation (25) to recover the known bounds in Sec. IV.
III IR kinematics and the acoustic metric
In order to study the propagation of the low-energy modes in this theory, it is useful to turn to the effective description given by the acoustic metric. The purpose of this Section is to identify IR kinematic quantities that are suitable to describe gapped modes, and that allow to synthetically express the known analyticity bounds. As we will discuss in Sec. IV, the usual notions of phase velocity and group/ray velocity, see e.g. [37, 38, 39], do not lead to a momentum-independent expression of the bounds. Instead, as we will see, a natural notion of velocity related to both phase and group velocity, and derived from the momentum of perturbations, encodes the known analyticity bounds through a simple, momentum-independent condition. We now proceed to summarize relevant information to be used in Sec. IV for discussing the IR characterization of the bounds.
Quite generally, the equations of motion for linearized wave-like excitations $\delta\chi\sim e^{ip_{\mu}x^{\mu}}$ of a field $\chi$ can be studied in terms of the excitation’s momentum $p_{\mu}$, obtaining an on-shell condition of the form:
$Z^{\mu\nu}(p)p_{\mu}p_{\nu}+C^{2}=0\,$, (27)
where $Z^{\mu\nu}$ is known as the acoustic metric, which we will introduce in detail below, see Sec. III.1, and $C^{2}$ is a constant that is associated with the gap of the $\delta\chi$ excitations.
In the case of an isotropic medium, this equation can be solved by a single dispersion relation for the components of $p_{\mu}=(-\omega,{\bf p})$: $\omega=\omega({\bf p})$. From the dispersion relation $\omega({\bf{p}})$, two a priori different notions of velocity are usually defined [^3] The appropriate definitions in the case of anisotropic media are discussed e.g. [40].:
$|{\bf v}_{p}|$ $=$ $\frac{\omega({\bf p})}{|{\bf p}|}\,,\quad\text{the phase velocity}\;,$ (28)
${\bf v}_{g}$ $=$ $\frac{\partial\omega({\bf p})}{\partial{\bf p}}\,,\quad\text{the group velocity}\;.$ (29)
The first one describes the propagation of constant phase surfaces of the wave, while the second is commonly associated with the motion of a wave packet. In the geometrical optics regime, the latter also corresponds to the velocity of particle-like trajectories/rays. Formally, the support in spacetime of the retarded Green’s function is ultimately given by the front/signal velocity, i.e. the limiting phase velocity at infinite momentum [23]  [^4] Though typically inaccessible from the low-energy EFT.. Both notions of velocity defined above can also be obtained from properly defined Lorentz-covariant velocity vectors $U_{p}^{\mu}$ and $U_{g}^{\mu}$ for the phase and group velocity respectively:
${\bf v}_{p}=\frac{{\bf U}_{p}}{U_{p}^{0}}\;,\quad{\bf v}_{g}=\frac{{\bf U}_{g}}{U_{g}^{0}}\;.$ (30)
The Lorentz-covariant phase velocity $U_{p}^{\mu}$ can be defined by
$U_{p}^{\mu}p_{\mu}=0$ (31)
up to a normalization, while the Lorentz-covariant group velocity $U_{g}^{\mu}$ is given by
$U_{g}^{\mu}=\frac{1}{2}\frac{\partial}{\partial p_{\mu}}\left(Z^{\alpha\beta}(p)p_{\alpha}p_{\beta}+C^{2}\right)\,.$ (32)
This is consistent with Eq. (30) because the equation of motion (EoM) implies
$\frac{\partial}{\partial{\bf p}_{i}}(Z^{\mu\nu}(p)p_{\mu}p_{\nu})+\frac{\partial}{\partial\omega}(Z^{\mu\nu}(p)p_{\mu}p_{\nu})\frac{\partial\omega({\bf p})}{\partial{\bf p}_{i}}=-\frac{d}{d{\bf p}_{i}}C^{2}=0\;,$ (33)
and we have defined $\omega=-p_{0}$.
By Eq. (27), we can recognize that $Z^{\mu\nu}$ defines the soundcone that characterizes on-shell excitations. For this reason, it is appropriate to treat $Z^{\mu\nu}$ as a metric for the momentum co-vector $p_{\mu}$, see also Eq. (39) and the related discussion. Then it is natural to define an acoustic vector in terms of $Z^{\mu\nu}$ and $p_{\mu}$:
$N^{\mu}=Z^{\mu\nu}(p)p_{\nu}\,$, (34)
which allows one to write the on-shell condition as
$N^{\mu}p_{\mu}=-C^{2}\,.$ (35)
The acoustic vector $N^{\mu}$ often coincides, up to a normalization, with either or both the phase velocity and the group velocity vectors. For instance, up to a normalization factor, we have
$\begin{split}N^{\mu}=&\,U_{p}^{\mu}\;,\qquad\text{for $C=0$}\;,\\ N^{\mu}=&\,U_{g}^{\mu}\;,\qquad\text{for $\frac{\partial Z^{\alpha\beta}}{\partial p_{\mu}}=0$}\;.\end{split}$ (36)
The three velocities coincide when there is no dispersivity in the system, i.e. when ${\bf v}_{p}$ and ${\bf v}_{g}$ are momentum independent: $N^{\mu}=U_{p}^{\mu}=U_{g}^{\mu}$. Conversely, in the general situation when neither of the conditions of Eq. (36) hold, all three velocity vectors are different $N^{\mu}\neq U_{p}^{\mu}\neq U_{g}^{\mu}$. Nevertheless, the acoustic vector $N^{\mu}$ still contains a valid notion of velocity:
${\bf v}$ $=$ $\frac{{\bf N}}{N^{0}}\,.$ (37)
In what follows we discuss the acoustic metric $Z^{\mu\nu}$ in more detail and motivate the use of the vector $N^{\mu}$ as a notion of velocity with which to express the known analyticity bounds. The reader interested only on the recovery of the bounds may skip the rest of this Section and instead go directly to Sec. IV. For a more detailed account of this formalism see Ref. [40].
III.1 Non-dispersive, gapless case
We now discuss the cases in which the system has second order equations for perturbations with no gap, and introduce the acoustic metric $Z^{\mu\nu}$ in detail. These assumptions amount to having $C=0$ and $\partial Z^{\alpha\beta}/\partial p_{\mu}=0$, meaning that the system has no dispersivity. For simplicity, we also consider an isotropic medium. We begin by considering a theory with gapless excitations $\delta\chi$ described by a quadratic Lagrangian of the form
@@ III.1 Non-dispersive, gapless case
where in this paragraph we assume $Z^{\mu\nu}$ to contain no differential operators, meaning that the theory has second order equations of motion. Here, for convenience, we have introduced $S_{\mu\nu}$ to be the inverse of $Z^{\mu\nu}$, and defined $S=\det(S_{\mu\nu})$. In this case the associated linearized EoM can be written as
@@
with $\bar{\square}=Z^{\mu\nu}\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}$ and $\bar{\nabla}_{\mu}$ is the covariant derivative metric-compatible with $Z$ and $S$. This is simply the wave-equation on the nontrivial background given by the acoustic metric.
In a general case in which $Z^{\mu\nu}(x)$ has a non-trivial space-time dependence, e.g. the background is not homogeneous, we study this equation in the Eikonal approximation, with an ansatz $\delta\chi=A(x)\,e^{i\mathcal{S}(x)/\epsilon}$, and expanding for $\epsilon\to 0$. This amounts to consider a rapidly varying phase $\mathcal{S}(x)$ and a slowly varying amplitude $A(x)$, appropriate for the geometric optics regime in which the wavelength of the excitation $\delta\chi$ is much shorter than the typical scales of the background inhomogeneities. In this limit, one can show that at leading order in $1/\epsilon$ Eq. (39) corresponds to the on-shell condition:
$Z^{\mu\nu}(x)p_{\mu}p_{\nu}=0\,$, (40)
where now $p_{\mu}=\partial_{\mu}\mathcal{S}$ is the momentum co-vector of the excitation. Since by assumption the theory is second order, $Z^{\mu\nu}$ does not itself depend on $p_{\mu}$.
We may construct a geometrical picture of how the excitations propagate in this regime by deriving a corresponding geodesic equation associated to $Z^{\mu\nu}$. Acting with Z-compatible vector derivative $\bar{\nabla}_{\alpha}$ on the above on-shell condition, we get
@@
where we used the fact that $p_{\mu}$ is the gradient of a scalar quantity, the phase $\mathcal{S}$, and therefore $\bar{\nabla}_{\alpha}p_{\mu}=\bar{\nabla}_{\mu}p_{\alpha}$. Finally, as in Eq. (34), we can define the acoustic vector $N^{\mu}$ by raising the index of the momentum co-vector $p_{\mu}$ with $Z^{\nu\alpha}$,
$N^{\mu}=Z^{\mu\nu}p_{\nu}\,.$ (42)
This leads to the following geodesic equation:
@@
In light of this, Eq. (41) can be rephrased as the parallel transport of energy-momentum along the $N^{\mu}$ geodesic, $N^{\mu}\bar{\nabla}_{\mu}p_{\nu}=0$. Physically, the acoustic vector $N^{\mu}$ tracks the flow of energy-momentum $p_{\mu}$. Note that, by definition of $N^{\mu}$, the EoM can be rewritten as
@@
In conclusion, the equations of motion in the gapless and dispersionless case imply that the acoustic vector $N^{\mu}$ follows null geodesics of $S_{\mu\nu}$. This defines the trajectories, also known as “rays”, describing propagation of the wavefronts in the geometric optics approximation:
$N^{\mu}p_{\mu}=N^{\mu}\partial_{\mu}\mathcal{S}=0\quad,\quad\text{for gapless, dispersionless modes.}$ (45)
This equation means that the vector $N^{\mu}$ defines surfaces of constant phase. In fact, as we discussed, in the present nondisperive cases $N^{\mu}$ coincides with the phase velocity as well as the group velocity vectors. Therefore, the corresponding velocity, Eq. (37),
${\bf v}=\frac{{\bf N}}{N^{0}}\,$,
coincides with both phase and group velocity, defined in Eqs. (28) and (29).
The description of kinematics in terms of the acoustic metric allows one to retrieve known results (see e.g. Appendix A), while enabling a geometrical interpretation of more complicated scenarios. As we will see, the velocity defined through the acoustic vector $N^{\mu}$ will remain a relevant quantity in dispersive cases, when the propagation is affected by higher derivatives as well as by a gap, and the equivalence between the different notions of velocity is no longer valid.
III.2 Gapped and dispersive systems
For our particular purpose of studying the gapped mode (25) in the low-momentum regime, we need to depart from the Eikonal approximation of the previous section and consider the regime in which momenta are not necessarily large enough to allow us to neglect the terms in the EoM with less derivatives. To keep the discussion simple, we consider a space-time independent acoustic metric and gap term. This is exactly the case for the model discussed in Sec. II, since there we have $\partial_{\rho}Z^{\mu\nu}=0$; see Eq. (6). This means that we can use a plane-wave ansatz $\delta\chi=A\,e^{ip_{\mu}x^{\mu}}$, with $A$ and $p_{\mu}$ constants. We thus obtain